OpenStudy (nerdsarecool):

Problem Page Multiply. Write your answer as a fraction in simplest form. ×578

OpenStudy (nerdsarecool):

Wait

OpenStudy (nerdsarecool):

It came out wrong

OpenStudy (anonymous):

There is some missing...I know it!

OpenStudy (nerdsarecool):

5 times 7/8

OpenStudy (nerdsarecool):

That is the question

OpenStudy (nerdsarecool):

5 times 7/8

OpenStudy (nerdsarecool):

4.37500 right

OpenStudy (anonymous):

Yes!

OpenStudy (nerdsarecool):

35/8

OpenStudy (nerdsarecool):

In a fraction in simplest form

OpenStudy (anonymous):

Yup.

OpenStudy (nerdsarecool):

Thank you again

OpenStudy (nerdsarecool):

I have two more

OpenStudy (anonymous):

Well, you got that one all on your own. :)

OpenStudy (nerdsarecool):

I can do it then

OpenStudy (anonymous):

Alright, I'll just double check if you like.

OpenStudy (nerdsarecool):

Ok

OpenStudy (nerdsarecool):

I am going to help some other people

OpenStudy (anonymous):

Ok.

OpenStudy (nerdsarecool):

0.818181818 as a fraction

OpenStudy (nerdsarecool):

What is that as a fraction

OpenStudy (anonymous):

Have any guesses?

OpenStudy (nerdsarecool):

Ummm

OpenStudy (nerdsarecool):

No

OpenStudy (nerdsarecool):

Sorry

OpenStudy (nerdsarecool):

Do you have any answers

OpenStudy (anonymous):

Do you have any choices?

OpenStudy (nerdsarecool):

No

OpenStudy (nerdsarecool):

Why is it to hard

OpenStudy (anonymous):

\[\frac{ 409090909 }{ 500000000 }\]

OpenStudy (nerdsarecool):

In simplest form

OpenStudy (nerdsarecool):

I will just do a different one

OpenStudy (anonymous):

Hang on.

OpenStudy (nerdsarecool):

Ok

OpenStudy (nerdsarecool):

Wi will

OpenStudy (nerdsarecool):

Hang on

OpenStudy (anonymous):

It is irreducible.

OpenStudy (nerdsarecool):

So no simplest form

OpenStudy (anonymous):

No more simple than it is.

OpenStudy (nerdsarecool):

It was wrong

OpenStudy (nerdsarecool):

I am going to leave

OpenStudy (anonymous):

Alright, sorry I couldn't help you better.

jimthompson5910 (jim_thompson5910):

hint: x = 0.8181818181..... 100x = 81.8181818.... 100x - x = 81.8181818.... - 0.8181818181..... 99x = ??? x = ???

jimthompson5910 (jim_thompson5910):

notice in the subtraction, the decimal portion `8181818....` gets eliminated

OpenStudy (anonymous):

So, could 9/11 be a possible solution?

jimthompson5910 (jim_thompson5910):

x = 0.8181818181..... 100x = 81.8181818.... 100x - x = 81.8181818.... - 0.8181818181.... 99x = 81 x = 81/99 x = 9/11 that is correct 9/11 = 0.81818181818181....

OpenStudy (anonymous):

Cool, thanks for the help in explaining it.

jimthompson5910 (jim_thompson5910):

no problem